OpenStudy (realslimshady):

Given the relation y = x2 + 1, identify the x- and y- axis intercepts.

OpenStudy (jbedu):

This will not be in real numbers but the value of x will be in complex number domain.

OpenStudy (jbedu):

Put x = 0, then y-intercept will be 1

OpenStudy (mathmale):

y = x2 + 1 would be clearer if written y = x^2 + 1. You need to use " ^ " to indicate exponentiation. At a y-intercept, x=0. To find the y-intercept, please let x=0 in y = x^2 + 1 and find the value of y. Write your y-intercept as (0, ? )

OpenStudy (jbedu):

But if you have y = 0, then x intercept will be x = \[\sqrt{-1}\]. So x - i (iota).

OpenStudy (jbedu):

x = i

OpenStudy (realslimshady):

this is trigonometry

OpenStudy (realslimshady):

actually it would be deez but thats besides the point

OpenStudy (mathmale):

Once again: "At a y-intercept, x=0. To find the y-intercept, please let x=0 in y = x^2 + 1 and find the value of y. Write your y-intercept as (0, ? )" Please note: y would NOT be 0 if x=0.

OpenStudy (realslimshady):

appreciate it